OpenStudy (rainbow_rocks03):

Will medal+fan for best response , no direct answers give the explanation. Determine the sum of the given probabilities. Are the events complementary? P(gold) = 4/11 ; P(silver) = 7/11

OpenStudy (rainbow_rocks03):

@BloomLocke367 @Ilovecake @Legends @LegendarySadist

OpenStudy (rainbow_rocks03):

combining in such a way as to enhance or emphasize the qualities of each other or another.

OpenStudy (rainbow_rocks03):

So I have to combine Gold & Silver , right?

OpenStudy (anonymous):

I suppose so, they wrote this one out really odd.

OpenStudy (rainbow_rocks03):

let me put the answer choices

OpenStudy (rainbow_rocks03):

A. 1 The events are complementary. B. 1 The events are not complementary. C. 1/2 The events are complementary. D. 1/2The events are not complementary.

OpenStudy (thomas5267):

Two events are complementary if and only if either of the event must occur, i.e. P(A)+P(B)=1.

OpenStudy (rainbow_rocks03):

I'm confused. :(

OpenStudy (rainbow_rocks03):

I don't understand what you wrote.

OpenStudy (thomas5267):

Put it this way. Two events are complementary if and only if the following is true: If A happens then B doesn't happen. If B happens then A doesn't happen. Either A or B must happen. Mathematically, P(A)+P(B)=1.

OpenStudy (rainbow_rocks03):

oh

OpenStudy (rainbow_rocks03):

so my answer is A. Is that correct?

OpenStudy (rainbow_rocks03):

@thomas5267

OpenStudy (thomas5267):

Yes.

OpenStudy (rainbow_rocks03):

thx for the help.

OpenStudy (thomas5267):

I should make this more clear. Two events are complementary if and only if ALL of the following are true: If A happens then B doesn't happen. If B happens then A doesn't happen. Either A or B must happen. Mathematically, P(A)+P(B)=1.